OpenStudy (anonymous):

could someone explain and answer this please? find the area of a square when a side is 3x + 5

OpenStudy (jmark):

Given 3x + 5 Area of the square = side * side Area of the square = (3x+5) (3x+5) Area of the square = (3x+5)^2 Area of the square = (5)^2+2(5)(3x)+(3x)^2 Area of the square = 25 - 30x + 9x^2 Follow us at : https://plus.google.com/+tutorvista/posts